# Scalable and shallow quantum circuits encoding probability distributions informed by asymptotic entanglement analysis

### v2

@@ A.3 Proof of Theorem 2
$\rho_{k}=\frac{1}{2^{k}}\sum_{m,n\geq 0}|f^{(m)}(u)\rangle\frac{1}{(1+m+n)n!m!2^{k(m+n)}}\langle f^{(n)}(u^{\prime})|,$ (34)
@@
$\rho_{k,m,n}=\sqrt{\langle f^{(n)}(u)|P_{n-1}|f^{(n)}(u)\rangle}\frac{1}{(m+n+1)m!n!2^{k(m+n)}}\sqrt{\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle},$ (36)
where $m,n\in\mathbb{N}_{0}$ and $\rho_{k,m,n}$ is a recomputed term under the sum in Eq.34. The scalar products $\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle$ appearing in the density matrix are defined for the function discretized at a finite resolution $1/2^{k}$, however, as before, via the Euler-MacLaurin formula they can be approximated by continuous integrals with boundary terms (which, as we work to leading order in $1/2^{k}$ throughout, we ignore). In particular, we have $\langle f^{(1)}(u)|P_{0}|f^{(1)}(u)\rangle=g_{1}(f)$.
@@
$e_{m}=(q_{0}/2^{mk},q_{1}/2^{(m-1)k},\dots,q_{m-1}/2^{k},1,q_{m+1}/2^{k},\dots)=(\vec{\tilde{q}},1,\vec{\tilde{Q}}),$ (37)
where we have introduced vectors $\vec{\tilde{q}},\vec{\tilde{Q}}$ with elements $\tilde{q}_{n}=q_{n}/2^{(m-n)k},\tilde{Q}_{n}=q_{m+1+n}/2^{k(n+1)}$. Now we need to solve the eigenvalue equation:
@@
where $\rho_{k,m}$ is the $m$-th eigenvalue of the $k$-th qubit reduced density matrix. To solve this equation one can write the matrix $\hat{\rho}_{k}$ in the block form:
@@
We begin solving the system from the second equation. Both $l\cdot\tilde{q}$ and $p$ are of the leading order $1/2^{2mk}$, while $w^{\dagger}\cdot\tilde{Q}=O(1/2^{(2m+2)k})$. Hence, to the leading order, we can safely neglect the term $w^{\dagger}\cdot\tilde{Q}$. Note also that for consistency $\rho_{k,m}=O(1/2^{2km})$. Consider now the $n$-th equation (with indexation starting from 00) from the matrix equation $M\tilde{q}+l+D\tilde{Q}=\rho_{k,m}\tilde{q}$. To the leading order we have $l_{n}=O(1/2^{k(m+n)})$, $M_{n}\tilde{q}=O(1/2^{k(m+n)})$, $D_{n}\tilde{Q}=O(1/2^{k(m+n+2)})$, $\rho_{k,m}\tilde{q}_{n}=O(1/2^{3km-nk})$ (note that $n<m$). Only first two terms remain and to the leading order we obtain:
@@
which allows us to find the vector $\tilde{q}$. Note that all the factors $1/2^{k}$ cancel out in these equations and we obtain a simple linear system on coefficients $q_{n}$ with $n<m$. $\rho_{k,m}$ can be found with substituting this solution for $\tilde{q}$ into $\rho_{k,m}=l^{\dagger}\cdot\tilde{q}+p$. Lastly, we should find the vector $\tilde{Q}$. Let us estimate the orders of elements of the $n$-th equation from $D^{\dagger}\tilde{q}+w+G\tilde{Q}=\rho_{k,m}\tilde{Q}$. We have $D^{\dagger}_{n}\tilde{q}=O(1/2^{k(2m+1+n)}$, $w_{n}=O(1/2^{k(2m+1+n})$, $\rho_{k,m}\tilde{Q}_{n}=O(1/2^{k(2m+1+n)})$, while $G_{n}\tilde{Q}=O(1/2^{k(2m+n+3)})$. Hence, we can neglect the term $G\tilde{Q}$ and just obtain $\rho_{k,m}\tilde{Q}=D^{\dagger}q+w$, where all $1/2^{k}$ factors also cancel out to the leading order.
@@
The matrix $N$ is the analog of the matrix $M$ for the eigenvalue $\rho_{k,m+1}$. To simplify the determinants we notice that the matrices $N,M$ can be factorized as follows: $N=F_{m+1}H_{m+1}F_{m+1},M=F_{m}H_{m}F_{m}$, where $F_{m}$ is a $m\times m$ diagonal matrix with diagonal elements $F_{m,nn}=\frac{\sqrt{\langle f^{(n)}(u)|P_{n-1}|f^{(n)}(u)\rangle}}{n!2^{kn}}$ ($n=0,1,\dots m-1$), while $H_{m}$ is a $m\times m$ matrix with matrix elements $H_{m,nl}=\frac{1}{n+l+1}$. With these factorizations we obtain the following formula for $\rho_{k,m}$:
$\rho_{k,m}=\left(\frac{\det{F_{m+1}}}{\det{F_{m}}}\right)^{2}\times\frac{\det{H_{m+1}}}{\det{H_{m}}}=\frac{\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle}{m!^{2}4^{km}}\times\frac{\det{H_{m+1}}}{\det{H_{m}}}.$ (46)
Here the first term contains information on the function and the $1/4^{km}$ scaling, while the second term $\frac{\det{H_{m+1}}}{\det{H_{m}}}$ is just a positive numerical coefficient, independent of $k$ and details of the function.
@@ B.2 Exact functional coefficients for probability distributions
Next we give the explicit expressions for the $g_{1}(f)$ functions for the symmetrical normal ($\mu=L/2$), log-normal and Lévy distributions, obtained by performing the integrals in Eq.30 for these examples:
$g_{1}(\sqrt{p(x)_{normal}})=\frac{L^{2}}{4\sigma^{2}}\left[1-\sqrt{\frac{L^{2}}{2\sigma^{2}}}\frac{\exp{\left(-\frac{L^{2}}{8\sigma^{2}}\right)}}{\left(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2})\right)}\right]$ (49)
$g_{1}(\sqrt{p(x)_{log-normal}})=\frac{L^{2}\exp{\left(-2\mu+2\sigma^{2}\right)}}{4\sigma^{2}\left[\sqrt{2\pi}-\Gamma\Bigl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma)^{2}}{2}\Bigr{)}/\sqrt{2}\right]}\times\\ \times\biggl{[}(1+\sigma^{2})\biggl{(}\sqrt{2\pi}-\Gamma\biggl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma+2\sigma)^{2}}{2}\biggr{)}/\sqrt{2}\biggr{)}+\\ +\left((2-\sqrt{2})\sigma-\frac{\log{(L\exp{(-\mu)})}}{\sqrt{2}\sigma}\right)\exp{[-(\log{(L\exp{[-\mu]})}/\sigma+2\sigma)^{2}/2]}\biggr{]}-\\ -\frac{L^{2}\exp{\{-2\mu+\sigma^{2}-(\log{(L\exp(-\mu))}/\sigma+\sigma)^{2}\}}}{4\sigma^{2}\times\left[\sqrt{2\pi}-\Gamma\Bigl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma)^{2}}{2}\Bigr{)}/\sqrt{2}\right]^{2}}$ (50)
$g_{1}(\sqrt{p(x)_{Levy}})=\frac{21L^{2}}{8c^{2}}+\frac{4L^{2}\exp{[-c/2L}]}{c^{2}\Gamma(1/2,c/2L)}\left[\frac{21}{16}\left(\frac{c}{2L}\right)^{1/2}+\frac{7}{8}\left(\frac{c}{2L}\right)^{3/2}+\right.\\ +\left.\frac{1}{8}\left(\frac{c}{2L}\right)^{5/2}+\frac{1}{4}\left(\frac{c}{2L}\right)^{7/2}\right]-\frac{L^{2}}{c^{2}\Gamma(1/2,c/2L)^{2}}\left(\frac{c}{2L}\right)^{3}\exp{[-c/L]}$ (51)
@@
$g_{2}(\sqrt{p(x)_{normal}})=\frac{L^{4}}{8\sigma^{4}}-\frac{L^{4}}{4\sigma^{4}}\left(\sqrt{\frac{L^{2}}{32\sigma^{2}}}+\sqrt{\frac{L^{2}}{8\sigma^{2}}}^{3}\right)\frac{\exp{\left[-\frac{L^{2}}{8\sigma^{2}}\right]}}{(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))}-\\ -\frac{L^{6}}{32\sigma^{6}}\frac{\exp{\left[-\frac{L^{2}}{4\sigma^{2}}\right]}}{(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))^{2}}$ (52)
If we drop the finitness of the interval, that is, consider a large $L$, the value becomes simpler:
$g_{2}(\sqrt{p(x)_{normal}})=1/8\sigma^{4}.$ (53)
@@ Appendix C Entanglement and function localization
In Sec.3.1 we showed that the entanglement on the smallest scales exhibits universal exponential decay. We can similarly ask how the entanglement behaves on the qubits corresponding to the largest scales.

### v3

@@ A.3 Proof of Theorem 2
$\rho_{k}=\frac{1}{2^{k}}\sum_{m,n\geq 0}\frac{|f^{(m)}(u)\rangle\langle f^{(n)}(u^{\prime})|}{(1+m+n)n!m!2^{k(m+n)}},$ (34)
@@
$\rho_{k,m,n}=$ (36)
$=\frac{\sqrt{\langle f^{(n)}(u)|P_{n-1}|f^{(n)}(u)\rangle\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle}}{(m+n+1)m!n!2^{k(m+n)}},$
where $m,n\in\mathbb{N}_{0}$ and $\rho_{k,m,n}$ is a recomputed term under the sum in Eq.34. The scalar products $\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle$ appearing in the density matrix are defined for the function discretized at a finite resolution $1/2^{k}$. However, as before, via the Euler-MacLaurin formula they can be approximated by continuous integrals with boundary terms (which, as we work to leading order in $1/2^{k}$ throughout, we ignore). In particular, we have $\langle f^{(1)}(u)|P_{0}|f^{(1)}(u)\rangle=g_{1}(f)$.
@@
$e_{m}$ $=(q_{0}/2^{mk},q_{1}/2^{(m-1)k},\dots,q_{m-1}/2^{k},1,q_{m+1}/2^{k},\dots)$
$=(\vec{\tilde{q}},1,\vec{\tilde{Q}}),$ (37)
where we have introduced vectors $\vec{\tilde{q}},\vec{\tilde{Q}}$ with elements $\tilde{q}_{n}=q_{n}/2^{(m-n)k},\tilde{Q}_{n}=q_{m+1+n}/2^{k(n+1)}$. We now need to solve the eigenvalue equation:
@@
where $\rho_{k,m}$ is the $m$-th eigenvalue of the $k$-th qubit reduced density matrix. To this end one can write the matrix $\hat{\rho}_{k}$ in the block form:
@@
We begin solving the system from the second equation. Both $l\cdot\tilde{q}$ and $p$ are of the leading order $1/2^{2mk}$, while $w^{\dagger}\cdot\tilde{Q}=O(1/2^{(2m+2)k})$. Hence, to the leading order, we can safely neglect the term $w^{\dagger}\cdot\tilde{Q}$. Note also that for consistency $\rho_{k,m}=O(1/2^{2km})$. Consider now the $n$-th equation (with indexation starting from 00) from the matrix equation $M\tilde{q}+l+D\tilde{Q}=\rho_{k,m}\tilde{q}$. To the leading order we have $l_{n}=O(1/2^{k(m+n)})$, $M_{n}\tilde{q}=O(1/2^{k(m+n)})$, $D_{n}\tilde{Q}=O(1/2^{k(m+n+2)})$, $\rho_{k,m}\tilde{q}_{n}=O(1/2^{3km-nk})$ (note that $n<m$). Only first two terms remain, and to the leading order we obtain:
@@
which allows to find the vector $\tilde{q}$. Note that all the factors $1/2^{k}$ cancel out in these equations and we obtain a simple linear system on coefficients $q_{n}$ with $n<m$. The eigenvalue $\rho_{k,m}$ can be found by substituting this solution for $\tilde{q}$ into $\rho_{k,m}=l^{\dagger}\cdot\tilde{q}+p$. Lastly, we should find the vector $\tilde{Q}$. Let us estimate the orders of elements of the $n$-th equation from $D^{\dagger}\tilde{q}+w+G\tilde{Q}=\rho_{k,m}\tilde{Q}$. We have $D^{\dagger}_{n}\tilde{q}=O(1/2^{k(2m+1+n)}$, $w_{n}=O(1/2^{k(2m+1+n})$, $\rho_{k,m}\tilde{Q}_{n}=O(1/2^{k(2m+1+n)})$, while $G_{n}\tilde{Q}=O(1/2^{k(2m+n+3)})$. Hence, we can neglect the term $G\tilde{Q}$ and just obtain $\rho_{k,m}\tilde{Q}=D^{\dagger}q+w$, where all $1/2^{k}$ factors also cancel out to the leading order.
@@
The matrix $N$ is the analog of the matrix $M$ for the eigenvalue $\rho_{k,m+1}$. To simplify the determinants we notice that the matrices $N,M$ can be factorized as follows: $N=F_{m+1}H_{m+1}F_{m+1},M=F_{m}H_{m}F_{m}$, where $F_{m}$ is a $m\times m$ diagonal matrix with diagonal elements given by:
$F_{m,nn}=\frac{\sqrt{\langle f^{(n)}(u)|P_{n-1}|f^{(n)}(u)\rangle}}{n!2^{kn}}$ (46)
for $n=0,1,\dots m-1$, while $H_{m}$ is a $m\times m$ matrix with matrix elements $H_{m,nl}=\frac{1}{n+l+1}$. With these factorizations we obtain the following formula for $\rho_{k,m}$:
$\rho_{k,m}$ $=\left(\frac{\det{F_{m+1}}}{\det{F_{m}}}\right)^{2}\times\frac{\det{H_{m+1}}}{\det{H_{m}}}$
$=\frac{\langle f^{(m)}(u)|P_{m-1}|f^{(m)}(u)\rangle}{m!^{2}4^{km}}\times\frac{\det{H_{m+1}}}{\det{H_{m}}}.$ (47)
Here the first factor contains information on the function and the $1/4^{km}$ scaling, while the factor $\frac{\det{H_{m+1}}}{\det{H_{m}}}$ is just a positive numerical coefficient, independent of $k$ and of the details of the function.
@@ B.2 Exact functional coefficients for probability distributions
Here we give explicit expressions for the $g_{1}(f)$ functions for the symmetrical normal ($\mu=L/2$), log-normal and Lévy distributions, obtained by performing the integrals in Eq.30 for these examples:
$g_{1}$ $(\sqrt{p(x)_{normal}})$ (50)
$=\frac{L^{2}}{4\sigma^{2}}\left[1-\sqrt{\frac{L^{2}}{2\sigma^{2}}}\frac{\exp{\left(-\frac{L^{2}}{8\sigma^{2}}\right)}}{\left(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2})\right)}\right]$
$g_{1}(\sqrt{p(x)_{log-normal}})=\\ =\frac{L^{2}\exp{\left(-2\mu+2\sigma^{2}\right)}}{4\sigma^{2}\left[\sqrt{2\pi}-\frac{1}{\sqrt{2}}\Gamma\Bigl{(}\frac{1}{2},\frac{((\log L-\mu)/\sigma)^{2}}{2}\Bigr{)}\right]}\times\\ \times\biggl{[}\frac{1+\sigma^{2}}{\sqrt{2}}\biggl{(}2\sqrt{\pi}-\Gamma\biggl{(}\frac{1}{2},\frac{((\log L-\mu)/\sigma+2\sigma)^{2}}{2}\biggr{)}\biggr{)}+\\ +\left((2-\sqrt{2})\sigma-\frac{\log L-\mu}{\sqrt{2}\sigma}\right)\exp{[\frac{(\frac{\log L-\mu}{\sigma}+2\sigma)^{2}}{-2}]}\biggr{]}-\\ -\frac{L^{2}\exp{\{-2\mu+\sigma^{2}-((\log L-\mu))/\sigma+\sigma)^{2}\}}}{2\sqrt{2}\sigma^{2}\left[2\sqrt{\pi}-\Gamma\Bigl{(}\frac{1}{2},\frac{((\log L-\mu)/\sigma)^{2}}{2}\Bigr{)}\right]^{2}}$ (51)
$g_{1}(\sqrt{p(x)_{Levy}})=\\ =\frac{21L^{2}}{8c^{2}}+\frac{4L^{2}\exp{[-c/2L}]}{c^{2}\Gamma(1/2,c/2L)}\left[\frac{21}{16}\left(\frac{c}{2L}\right)^{1/2}+\frac{7}{8}\left(\frac{c}{2L}\right)^{3/2}+\right.\\ +\left.\frac{1}{8}\left(\frac{c}{2L}\right)^{5/2}+\frac{1}{4}\left(\frac{c}{2L}\right)^{7/2}\right]+\\ -\frac{L^{2}}{c^{2}\Gamma(1/2,c/2L)^{2}}\left(\frac{c}{2L}\right)^{3}\exp{[-c/L]}$ (52)
@@
$g_{2}(\sqrt{p(x)_{normal}})=\\ =\frac{L^{4}}{8\sigma^{4}}-\frac{L^{4}}{4\sigma^{4}}\left(\sqrt{\frac{L^{2}}{32\sigma^{2}}}+\sqrt{\frac{L^{2}}{8\sigma^{2}}}^{3}\right)\frac{\exp{\left[-\frac{L^{2}}{8\sigma^{2}}\right]}}{\Gamma(\frac{1}{2})-\Gamma(\frac{1}{2},\frac{L^{2}}{8\sigma^{2}})}-\\ -\frac{L^{6}}{32\sigma^{6}}\frac{\exp{\left[-\frac{L^{2}}{4\sigma^{2}}\right]}}{[\Gamma(\frac{1}{2})-\Gamma(\frac{1}{2},\frac{L^{2}}{8\sigma^{2}})]^{2}}$ (53)
If we drop the finitness of the interval, that is, consider the large $L$ limit, the expression simplifies to:
$g_{2}(\sqrt{p(x)_{normal}})=\frac{1}{8\sigma^{4}}.$ (54)
@@ Appendix C Entanglement and function localization
In Sec.III.1 we showed that the entanglement on the smallest scales exhibits universal exponential decay. We can similarly ask how the entanglement behaves on the qubits corresponding to the largest scales.